HIGH RESOLUTION FT STUDY OF THE INFRARED AND VISIBLE EMISSION SPECTRA OF AgH AND AgD

\underline{DOMINIQUE R.T. APPADOO}, K. ANDERSON, I. GORDON, A. SHAYESTEH, AND P.F. BERNATH, UNIVERSITY OF WATERLOO, 200 UNIVERSITY AVENUE, WATERLOO, ON N2L 3G1, CANADA.

The infrared and visible (A-X) emission spectra of^{107/109}AgH and^{107/109}AgD have been recorded with a Fourier transform spectrometer. The silver hydride molecules were generated by heating Ag metal to 1050 − 1325^{o}C in a high-tempeature furnace containing a few torr of an Ar/H mixture subjected to a 3kV/330mA discharge. The observed spectra consist mainly of rovibronic transitions belonging to the Δv = 0,_{−}^{+} 1,_{−}^{+} 2 sequences. The higher quality measurements from this study together with previous grating data on the A-X system of AgH, and micrawave data were first treated in a combined isotopomer fit to yield improved empirical molecular parameters for both the ground and first excited electronic states of each species. The A-state potential curve of AgH has an anomalous shape resulting from avoided crossings. In an attempt to better determine the A-state potential curve and its mass-dependent Born-Oppenheimer breakdown parameters, the combined data were directly fitted to potential energy functions.